Minutes of the 19th SOHO SWT Meeting

Goddard Space Flight Center, Greenbelt, MD 16 May 1996 Contents

1 Action Items SOHO SWT 19 2 2 Agree Agenda and Actions Revision 2 3 S/C Status 2 4 Ground System Status 2 5 Status of Experiments 2 6 Experiment Anomaly Reporting 3 7 Future Roll Maneuvers 3 8 Data Issues 3 9 Priorities for the Coming Three Months 3 10 SOL’s 4 11 PR Matters 4 12 SOHO Science Report 4 13 Guest Investigator Programme 4 14 Solar Physics Special: First Results from SOHO 5 15 SOHO Workshops and Colloquia 5 16 AOB 5 Annex 1: Agenda 6 Annex 2: List of Participants 8 Annex 3: Spacecraft Status 10 Annex 4: Ground System Status 24 Annex 5: Status of Experiments 34 Annex 6: Data Issues 45

1 Action Items SOHO SWT 19

Action 19-1: On MDI, SUMER, EIT/LASCO, UVCS to provide each month solar center coordinates for assessment of medium term pointing stability. Input to PS by 20th of each month.

Action 19-2: On PS to request reprocessing effort of all level 0 data from tapes. Due immediately.

Action 19-3: On PIs to revisit number of CD-ROM copies they need. Input to PS before 15 June.

2 Agree Agenda and Actions Revision

Agenda — annex 1 List of Participants — annex 2

Actions revision

Action 18-1: closed on PIs to check data quality and completeness of CD-ROMs. Input to PS by 28 February.

Action 18-2: obsolete on PIs to provide additional names for distribution list of PR package which will be circulated prior to the 2 May media event. Input to PS before 1 April.

Action 18-3: open on PIs to provide input for the next SOHO Science Report. Input to PS before 15 March.

3 S/C Status

H. Schweitzer gave a status report on the power subsystem (see annex 3). M. Hervieux presented commissioning and trending analysis results and explained the present strategy for thermoelastic operations (see annex 3). Several of the instruments, in particular VIRGO, GOLF and MDI, are very sensitive to temperature variations. After some discussion it was agreed to have a splinter meeting to work out a new strategy for medium term pointing adjustments without changing the S/C temperature. For outcome of this splinter see Minutes of 17 May SPWG meeting, and Vicente Domingo’s memo on mid-term pointing stability from 4 June 1996. There is an action on MDI, SUMER, EIT/LASCO, UVCS to provide each month solar center coordinates for assessment of medium term pointing stability. Input to PS by 20th of each month (Action 19-1).

4 Ground System Status

E. Larduinat gave a presentation on the status of the ground system (see annex 4). New limits: no input received yet from CEPAC and LASCO. POCC crashes were not considered to have big impact on operations.

5 Status of Experiments

– see annex 5

6 Experiment Anomaly Reporting

It was agreed to continue with the anomaly reporting on the Web (http://orpheus.nascom.nasa.gov/cgi-bin/anomaly report and http://orpheus.nascom.nasa.gov/cgi-bin/ anomaly browse) and add a special entry for S/C anomalies. Access to the SOHO Anomaly Reporting System is strictly controlled and is open only to certain workstations in a list of authorized hosts. If you need access, please contact L. Sanchez, D. Zarro, or Liyun Wang.

7 Future Roll Maneuvers

It was agreed to go ahead with the 90^{∘} roll on May 20/21, with short stops at 45^{∘} on the way to and back from 90^{∘}. Decision on which pole, north or south, to be taken at weekly planning meeting tomorrow. Next SWT meeting will decide when next roll will occur.

8 Data Issues

V. Buczkowski gave an overview of the level 0 data processing (see annex 6). There is a delay of approximately 4 to 5 days between the real time data reception and the 8 mm tapes from DSN. V. Buczkowski will make information available on Web on when CD-ROMs were sent to experimenters and with what data. There is an action on the Project Scientist to request reprocessing effort of all level 0 data from 8 mm tapes (Action 19-2). There is an action on the PIs to revisit number of CD-ROM copies they need. Input to PS before 15 June (Action 19-3). L. Sanchez gave an overview of the level-0 data distribution, data quality, status of summary data, and SOHO archives (see annex 6).

9 Priorities for the Coming Three Months

JOPs run so far:

• Intercal001 (CDS, SUMER, EIT)

• Intercal002, phase 2a and 2b (CDS + SEM)

• Intercal007 Stellar crosscalibration

• Intercal009 Above limb calibration

• JOP 002 Temperature Gradient in a CH (test; no rolls)

• JOP 012 Limb prominences (joint obs. with Pic du Midi)

• JOP 013 Temporal variability in the Transition Region

• JOP 015 Explosive Events

• JOP 022 Network Study

• JOP 023 XBPs (campaign with Yohkoh and GBOs)

• JOP 028 Streamers

• JOP 035 Coronal Hole Structure and Evolution

• JOP 039 Polar Plume Flow Priorities for the coming 3 months unchanged, i.e.:

1. Coronal Holes

2. Streamers

3. Diffuse Corona

4. Quiet Network

JOP leaders have to be present at the EOF when their JOP is run for the first time. To optimize JOPs, a report is due before re-run. Proposed campaigns and collaborations: kept up-to-date on Web (see monthly calender at http://sohowww.nascom.nasa.gov/operations)

10 SOL’s

It was agreed to implement the Science Operation Leader (SOL) concept as described in the SOP in June. Each week a different team is reponsible for providing a SOL, including the non-resident PI teams.

11 PR Matters

V. Domingo briefly reviewed the 2 May event which was considered a great success. There was a wide media coverage both in Europe as well as in the US, including daily newspapers, magazines and TV. H. Kunow had three points which should be improved next time: a) the PIs were not properly introduced and therefore had no visibility to the press; b) it was too stiff, and c) the press release did not mention the in-situ experiments.

The project scientist team is working on a “SOHO Image Gallery” on the Web showing sample images/data of all experiments. Starting point will be the data presented at the 2 May event in Paris. Awaiting DAT tapes from 4:2:2 videographics in Bristol.

It is planned to have future press releases in fairly regular intervals of 2 to 3 months, and whenever something spectacular has been observed or discovered. Some of these releases will be simple press releases, some in a format similar to the Space Science Updates of NASA PA. In addition, there should be SOHO press releases at the major meetings (AAS, COSPAR, etc.).

12 SOHO Science Report

Still awaiting input from experimenters (see action 18-3).

13 Guest Investigator Programme

A SOHO Guest Investigator (GI) programme has been announced by ESA and NASA on May 20 (ESA SCI(96)6 and NRA 96-OSS-09). Details of the programme can be found on the Internet at http://sohowww.nascom.nasa.gov/operations/guest investigators. The schedule of the GI programme is:

20 May 1996 Issue of AO 20 July 1996 Letter of Intent due 20 September 1996 Proposals due 20 December 1996 GI Selection announced 20 March 1997 Star of GI investigation 19 March 1997 Enf of investigation period 20 April 1998 Final report due

14 Solar Physics Special: First Results from SOHO

• flexible approach: split in 3 parts (Nov 96, Jan 97, March 97)

• Schedule:

– 15 June deadline for submission of papers

– 15 July papers to printer

– 1 August proofs

– 1 September final papers/proofs to printer

• meeting 15 June deadline is no guarantee for November publication

• Papers to be submitted directly to:

Kluwer Academic Publishers Editorial Office Attn: Mrs Sofie van der Waaij P.O. Box 17 3300 AA Dordrecht The Netherland,

clearly indicating that it is for SOLAR PHYSICS, First Results of SOHO

15 SOHO Workshops and Colloquia

The 1997 ESLAB Symposium will be devoted to the STSP cornerstone. Potential dates are 9-13 June or 6-10 October. There are also plans to organize a 5th SOHO Workshop in the first half of 1996, in continuation of the series of SOHO workshops held before launch.

16 AOB

Next SWT meeting: 15/16 August 1996 (again in conjunction with the monthly SPWG meeting)

Annex 1: Agenda

6

Annex 2: List of Participants

8

Annex 3: Spacecraft Status

10

Annex 4: Ground System Status

24

Annex 5: Status of Experiments

34

Annex 6: Data Issues

45